53% of all violent felons in the prison system are repeat offenders. if 43 violent felons are randomly selected, find the probability that..
a. Exactly 24 of them are repeat offenders
b. At most 25 of them are repeat offenders
c. At least 24 of them are repeat offenders
d. between 17 and 25 (including 17 and 25) of them are repeat offenders
Answer #1

a)

p = 0.53

n =43

P ( X =    24   ) = C(43,24) * 0.53^24 * (1-0.53)^19 =            0.1137

b)

X P(X)
P ( X = 0) = C (43,0) * 0.53^0 * ( 1 - 0.53)^43= 0 0.0000
P ( X = 1) = C (43,1) * 0.53^1 * ( 1 - 0.53)^42= 1 0.0000
P ( X = 2) = C (43,2) * 0.53^2 * ( 1 - 0.53)^41= 2 0.0000
P ( X = 3) = C (43,3) * 0.53^3 * ( 1 - 0.53)^40= 3 0.0000
P ( X = 4) = C (43,4) * 0.53^4 * ( 1 - 0.53)^39= 4 0.0000
P ( X = 5) = C (43,5) * 0.53^5 * ( 1 - 0.53)^38= 5 0.0000
P ( X = 6) = C (43,6) * 0.53^6 * ( 1 - 0.53)^37= 6 0.0000
P ( X = 7) = C (43,7) * 0.53^7 * ( 1 - 0.53)^36= 7 0.0000
P ( X = 8) = C (43,8) * 0.53^8 * ( 1 - 0.53)^35= 8 0.0000
P ( X = 9) = C (43,9) * 0.53^9 * ( 1 - 0.53)^34= 9 0.0000
P ( X = 10) = C (43,10) * 0.53^10 * ( 1 - 0.53)^33= 10 0.0001
P ( X = 11) = C (43,11) * 0.53^11 * ( 1 - 0.53)^32= 11 0.0002
P ( X = 12) = C (43,12) * 0.53^12 * ( 1 - 0.53)^31= 12 0.0005
P ( X = 13) = C (43,13) * 0.53^13 * ( 1 - 0.53)^30= 13 0.0014
P ( X = 14) = C (43,14) * 0.53^14 * ( 1 - 0.53)^29= 14 0.0033
P ( X = 15) = C (43,15) * 0.53^15 * ( 1 - 0.53)^28= 15 0.0073
P ( X = 16) = C (43,16) * 0.53^16 * ( 1 - 0.53)^27= 16 0.0144
P ( X = 17) = C (43,17) * 0.53^17 * ( 1 - 0.53)^26= 17 0.0258
P ( X = 18) = C (43,18) * 0.53^18 * ( 1 - 0.53)^25= 18 0.0420
P ( X = 19) = C (43,19) * 0.53^19 * ( 1 - 0.53)^24= 19 0.0624
P ( X = 20) = C (43,20) * 0.53^20 * ( 1 - 0.53)^23= 20 0.0844
P ( X = 21) = C (43,21) * 0.53^21 * ( 1 - 0.53)^22= 21 0.1042
P ( X = 22) = C (43,22) * 0.53^22 * ( 1 - 0.53)^21= 22 0.1175
P ( X = 23) = C (43,23) * 0.53^23 * ( 1 - 0.53)^20= 23 0.1210
P ( X = 24) = C (43,24) * 0.53^24 * ( 1 - 0.53)^19= 24 0.1137
P ( X = 25) = C (43,25) * 0.53^25 * ( 1 - 0.53)^18= 25 0.0975

= 0.7957

c)

X P(X)
P ( X = 0) = C (43,0) * 0.53^0 * ( 1 - 0.53)^43= 0 0.0000
P ( X = 1) = C (43,1) * 0.53^1 * ( 1 - 0.53)^42= 1 0.0000
P ( X = 2) = C (43,2) * 0.53^2 * ( 1 - 0.53)^41= 2 0.0000
P ( X = 3) = C (43,3) * 0.53^3 * ( 1 - 0.53)^40= 3 0.0000
P ( X = 4) = C (43,4) * 0.53^4 * ( 1 - 0.53)^39= 4 0.0000
P ( X = 5) = C (43,5) * 0.53^5 * ( 1 - 0.53)^38= 5 0.0000
P ( X = 6) = C (43,6) * 0.53^6 * ( 1 - 0.53)^37= 6 0.0000
P ( X = 7) = C (43,7) * 0.53^7 * ( 1 - 0.53)^36= 7 0.0000
P ( X = 8) = C (43,8) * 0.53^8 * ( 1 - 0.53)^35= 8 0.0000
P ( X = 9) = C (43,9) * 0.53^9 * ( 1 - 0.53)^34= 9 0.0000
P ( X = 10) = C (43,10) * 0.53^10 * ( 1 - 0.53)^33= 10 0.0001
P ( X = 11) = C (43,11) * 0.53^11 * ( 1 - 0.53)^32= 11 0.0002
P ( X = 12) = C (43,12) * 0.53^12 * ( 1 - 0.53)^31= 12 0.0005
P ( X = 13) = C (43,13) * 0.53^13 * ( 1 - 0.53)^30= 13 0.0014
P ( X = 14) = C (43,14) * 0.53^14 * ( 1 - 0.53)^29= 14 0.0033
P ( X = 15) = C (43,15) * 0.53^15 * ( 1 - 0.53)^28= 15 0.0073
P ( X = 16) = C (43,16) * 0.53^16 * ( 1 - 0.53)^27= 16 0.0144
P ( X = 17) = C (43,17) * 0.53^17 * ( 1 - 0.53)^26= 17 0.0258
P ( X = 18) = C (43,18) * 0.53^18 * ( 1 - 0.53)^25= 18 0.0420
P ( X = 19) = C (43,19) * 0.53^19 * ( 1 - 0.53)^24= 19 0.0624
P ( X = 20) = C (43,20) * 0.53^20 * ( 1 - 0.53)^23= 20 0.0844
P ( X = 21) = C (43,21) * 0.53^21 * ( 1 - 0.53)^22= 21 0.1042
P ( X = 22) = C (43,22) * 0.53^22 * ( 1 - 0.53)^21= 22 0.1175
P ( X = 23) = C (43,23) * 0.53^23 * ( 1 - 0.53)^20= 23 0.1210

= 1 - 0.5846

= 0.4154

d)

P ( X = 17) = C (43,17) * 0.53^17 * ( 1 - 0.53)^26= 17 0.0258
P ( X = 18) = C (43,18) * 0.53^18 * ( 1 - 0.53)^25= 18 0.0420
P ( X = 19) = C (43,19) * 0.53^19 * ( 1 - 0.53)^24= 19 0.0624
P ( X = 20) = C (43,20) * 0.53^20 * ( 1 - 0.53)^23= 20 0.0844
P ( X = 21) = C (43,21) * 0.53^21 * ( 1 - 0.53)^22= 21 0.1042
P ( X = 22) = C (43,22) * 0.53^22 * ( 1 - 0.53)^21= 22 0.1175
P ( X = 23) = C (43,23) * 0.53^23 * ( 1 - 0.53)^20= 23 0.1210
P ( X = 24) = C (43,24) * 0.53^24 * ( 1 - 0.53)^19= 24 0.1137
P ( X = 25) = C (43,25) * 0.53^25 * ( 1 - 0.53)^18= 25 0.0975

= 0.7685

Please revert back in case of any doubt.

Please upvote. Thanks in advance.